I have 4 spinner and there are 3 colours on the spinner red, blue and orange. If i spin the spinner twice what is the probability that i get all the same colour on the first spin and all the same colour on the second spin.
Found 2 solutions by venugopalramana, longjonsilver:
Answer by venugopalramana(3286)

I have 4 spinner
WHAT IS A 4 SPINNER?
and there are 3 colours on the spinner red, blue and orange. If i spin the spinner twice what is the probability that i get all the same colour
WHAT DOES THIS MEAN...PLEASE EXPLAIN .IN ONE SPIN YOU GET 1 COLOUR COMING UP..WHAT IS MEANT BY ALL THE SAME COLOUR
on the first spin and all the same colour on the second spin.
SAME QUESTION
IF YOU MEAN ONE COLOUR COMING UP ON I SPIN AND SAME COLOUR COMING UP ON II SPIN ,I CAN UNDERSTAND.
THAT PROBABILITY IS..ASSUMING 4 SPINNER HAS 3 COLOURS AND ONE BLANK...
ANY COLOUR ON I SPIN =3/4
SAME COLOUR ON II SPIN =1/4
PROBABILITY = (3/4)*(1/4)=3/16

Answer by longjonsilver(2297)

Assuming all the spinners are unbiased and the same and that P(red)=P(blue)=P(orange), then we have:

P(rad) = P(blue) = P(orange) = 1/3

On first spin of the 4 tops:
P(all red) = (1/3) * (1/3) * (1/3) * (1/3)
P(all red) = (1/81)

Similarly, P(all blue) = (1/81) and P(all orange) = (1/81)

So, P(all 4 tops are the same colour) = (1/81) + (1/81) + (1/81)
P(all 4 tops are the same colour) = (3/81)
P(all 4 tops are the same colour) = 1/27

Now, spinning the top a second time produces the same result... the 2 situations are independent, since they do not affect each other.

So, P(4 same colour AND then 4 same colour again) = (1/27) * (1/27)
--> 1/729

jon.